OpenStudy (anonymous):

i need help with area of a circle A circular plate has a diameter of 6 inches. What is the exact area of the plate? A.3π square inches B.6π square inches C.36π square inches D.9π square inches

OpenStudy (anonymous):

What's the formula for area of a circle?

OpenStudy (anonymous):

i don't know that is y i need help

OpenStudy (anonymous):

Alright so, Area of a circle = pir^2

OpenStudy (anonymous):

\[A=\pi r^2\]

OpenStudy (anonymous):

So since you have diameter, you know the radius is half of the diameter.

OpenStudy (anonymous):

Or you can say it as: \[A=\frac{ 1 }{ 4 }\pi d^2\]

OpenStudy (anonymous):

Either way you'll get the same answer so you can substitude A = pi*3^2 or use the formula above with 6 :)

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

So what's your answer?

OpenStudy (anonymous):

I wrote it :P, just point it out for me.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

What's your final answer? From your choices, with the information I gave you, which one will it be, I even wrote the answer.

OpenStudy (anonymous):

i think it is d

OpenStudy (anonymous):

Nope, remember I said radius = half of the diameter. And the area of a circle = pi*r^2

OpenStudy (anonymous):

Half of 6 = 3, so your answer is A.

OpenStudy (anonymous):

Oh wait

OpenStudy (anonymous):

it is c

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

You were right, it's D because 3^2 = 9

OpenStudy (anonymous):

o ok hahaha

OpenStudy (anonymous):

I didn't see the units :P

OpenStudy (anonymous):

9(pi) in^2 the units haha, that's where they get you

OpenStudy (anonymous):

ya